### v1

# Condor: A Code Discriminator Integrating General Semantics with Code Details

Abstract.
LLMs demonstrate significant potential across various software engineering tasks. However, they still face challenges in generating correct code on the first attempt when addressing complex requirements. Introducing a discriminator to select reliable outputs from multiple generated results is an effective way to enhance their reliability and stability. Currently, these discriminators fall into two categories: execution-based discriminators and non-execution-based discriminators. Execution-based discriminators face flexibility challenges due to difficulties in obtaining test cases and security concerns, while non-execution-based discriminators, although more flexible, struggle to capture subtle differences in code details. To maintain flexibility while improving the model’s ability to capture fine-grained code details, this paper proposes Condor. We first design contrastive learning to optimize the code representations of the base model, enabling it to reflect differences in code details. Then, we leverage intermediate data from the code modification process to further enrich the discriminator’s training data, enhancing its ability to discern code details. Experimental results indicate that on the subtle code difference dataset (i.e., CodeNanoFix), Condor significantly outperforms other discriminators in discriminative performance: Condor (1.3B) improves the discriminative F1 score of DeepSeek-Coder (1.3B) from 67% to 73%. In discriminating LLM-generated outputs, Condor (1.3B) and Condor (110M) raise the Pass@1 score of Meta-Llama-3.1-Instruct (70B) on the CodeNanoFix dataset from 52.64% to 62.63% and 59.64%, respectively. Moreover, Condor demonstrates strong generalization capabilities on the MBPP and APPS datasets. For example, Condor (1.3B) improves the Pass@1 of Meta-Llama-3.1-Instruct (70B) on the APPS dataset by 147.05%.
Code Discriminator, Code Embedding, Code Generation, Code Repair
copyright: noneccs: Software and its engineering Automatic programming
1. Introduction
With the widespread application of large language models (LLMs), their potential for tasks in software engineering (e.g., code generation and code repair) receives significant attention (Guo et al., 2024; Dubey et al., 2024; Team, 2024; Nam et al., 2024; OpenAI, 2024). However, they typically struggle to produce correct code on the first attempt when addressing complex requirements such as advanced algorithmic problems (Hendrycks et al., 2021; Yetiştiren et al., 2023; Khan et al., 2023; Ji et al., 2023). Existing approaches typically employ an automated discriminator to select a correct result from multiple outputs generated by LLMs, improving the reliability and stability of LLMs when handling complex problems (Inala et al., 2022).
Currently, these approaches for code discrimination fall into two main categories: execution-based approaches and non-execution-based approaches. Execution-based approaches require running the code before making a discriminative decision, utilizing feedback from execution as additional detailed information to enhance the discriminator’s performance. However, execution-based models face challenges in flexibility. First, executing code poses challenges because (1) many real-world downstream reasoning tasks lack test cases (Briand, 2007), (2) LLM-generated code may contain incomplete snippets (Chen et al., 2021a), and (3) the testing environments and other dependencies are not always accessible (Inala et al., 2022; cop, 2024). Second, even when execution is possible, running LLM-generated code raises safety concerns. On the other hand, non-execution-based approaches do not execute code and rely solely on the details presented in the code text, providing greater flexibility and applicability across different application scenarios.
However, existing non-execution-based discriminators typically leverage general code representation capabilities without adapting to the nuanced details of the programming process (Inala et al., 2022; Zhang et al., 2023). Challenges remain in improving the discriminator’s effectiveness at capturing intricate code details. In the development scenario, the functionality of code often depends on key minor changes, such as variable updates, conditional logic modifications, or adjustments in function calls. For example, as shown in Figure 1, the upper section illustrates a problem description about finding a number K equally distant from two distinct numbers A and B. Submission 1 displays buggy code, while submission 2 shows the correct code after the user’s modification. Notably, only three characters differ between the two submissions (i.e., ‘(’, ‘)’, and ‘/’), yet the functionality of the first submission remains incorrect. Similarly, outputs generated by LLM can also contain errors caused by subtle differences, demonstrating that such mistakes are not unique to humans. For example, as shown below Figure 1, the model (i.e., Meta-LLama-3.1-Instruct-70B) generates two outputs for Task 173 in the MBPP (Austin et al., 2021) benchmark, one of which is incorrect due to a minor substitution, where ‘all’ is mistakenly replaced with ‘any’. Existing discriminators primarily rely on embeddings provided by other general-purpose code models as a basis for representation. This limits their ability to capture subtle code details, consequently reducing the effectiveness of identifying similar sample pairs in LLM-generated outputs.
Figure 1. An example illustrates how humans write correct code through thinking, and how models rely on a discriminator to select the correct code. The upper section illustrates the interaction with the code evaluation system, where the user attempts to submit the code twice. The lower section displays the correct code selection by a discriminator, where the model may not be able to generate the correct answer on its first attempt. Thus, it is necessary to employ a discriminator to enhance the reliability of the generated outputs.
To keep the flexibility and enhance the detail-capturing capabilities of the code discriminator, we propose Condor, a non-execution-based discriminator that integrates general semantics with code details. To address the challenges that non-execution-based approaches face in capturing subtle code differences, we design an embedding-level contrastive learning strategy (Chen et al., 2020) and a data-level intermediate data mining strategy for Condor. In our embedding-level contrastive learning strategy, we first sample code pairs with high textual similarity for the same problem. Then, we design two loss functions during the training process to progressively bring the embeddings of correct code samples closer together, while increasing the embedding distance between correct and incorrect code snippets. Condor refines the base model’s code embeddings through contrastive learning to better distinguish between correct and incorrect code representations. For the data-level intermediate data mining strategy, we utilize data from multiple submissions for the same problem to identify intermediate data not covered by the existing training set. This strategy aims to capture intermediate data that the evaluation system may not have recorded during the user’s problem-solving process. For example, in the upper section of Figure 1, after the user’s thinking process (i.e., h4. Thinking), the user submits a new code snippet (i.e., submission 2) with modifications in two steps. After the modification in Step 1, the code still contains a bug. This partially modified version, which has not yet undergone the Step 2 modification, is referred to as partially fixed intermediate data. In real-world scenarios, users may submit a new code version only after careful consideration, and the partially fixed code for previous submissions may not always be explicitly recorded in the evaluation system. By leveraging these partially fixed buggy code snippets, Condor can incorporate the incremental details in the programmer’s thinking process to further enhance its discriminative capabilities in specific tasks. As shown in the lower section of Figure 1, by integrating these two strategies for capturing code details, Condor aims to automatically select the correct code from multiple outputs when the LLM struggles to generate the correct code on the first attempt.
We propose two parameter-scale versions of Condor (i.e., 110M and 1.3B) and conduct extensive experiments to evaluate its effectiveness. We find that current datasets are insufficient for fully quantifying the discriminator’s ability to discern differences in code details (Chen et al., 2021b; Muennighoff et al., 2023). Thus, we construct the CodeNanoFix dataset, which includes a series of samples consisting of an algorithmic problem description, pairs of the code with subtle errors, and the corresponding correct version. First, we apply the Condor to the classification task to assess the model’s ability to recognize subtle code differences. The results show that general code generation models struggle to accurately distinguish these fine-grained differences, often resulting in imbalances in precision or recall, leading to suboptimal overall performance. In contrast, Condor excels across precision, recall, and F1-score, significantly outperforming other models. For example, Condor (110M) improves the F1-score from around 67% to 72.34%. Additionally, Condor (1.3B) demonstrates substantial improvements in discriminative capability when selecting the correct code from multiple generated outputs, achieving an increase in Pass@1 accuracy of Meta-Llama-3.1-Instruct (70B) from 52.64% to 62.63%. Second, to validate Condor’s generalization ability, we test it on the APPS (Hendrycks et al., 2021) and the MBPP (Austin et al., 2021) datasets. The results indicate that Condor consistently maintains strong performance across different datasets. For example, on the APPS dataset, Condor (1.3B) raises the Pass@1 of Meta-Llama-3.1 Instruct (70B) from 10.16% to 25.10%, and on MBPP, Condor (1.3B) improves Pass@1 from 71.40% to 75.20%. Finally, through ablation studies, we demonstrate that the two strategies for capturing code details in Condor are crucial to its overall performance, further validating their effectiveness in detecting subtle code differences.
We summarize our contributions as follows:
  • We propose Condor, a novel, non-execution-based code discriminator that incorporates embedding-level and data-level code details to discriminate and filter code generated by LLMs.
  • We construct CodeNanoFix, a new code dataset focusing on bugs caused by subtle code changes, to evaluate whether the discriminators can distinguish fine-grained differences in code.
  • We conduct extensive experiments, and results show that Condor effectively differentiates between code that appears similar in form but differs in functionality.
2. Condor
2.1. Overview of Condor
Figure 2. The Condor overview consists of two main components: contrastive learning at the embedding level to capture code details (upper section), and data-level augmentation through intermediate code, which supplements code details that are not recorded in existing datasets (lower section).
Condor is a model designed to discriminate the correctness of code purely based on its text, without the need to execute the code to gather additional information. This gives Condor greater flexibility in real-world applications, especially when test cases are unavailable.
However, distinguishing between code samples that are very similar in text becomes more challenging for the discriminator, as it may struggle to fully utilize code details to differentiate between code quality or functionality. To address this challenge, Condor, as illustrated in Figure 2, employs two strategies to enhance its ability to capture fine-grained code details.
  • Embedding-Level Contrast on Code Details: We design a contrastive learning strategy to capture code details at the embedding level.
  • Data-Level Augmentation on Code Details: We utilize intermediate code details, which may be missing from existing datasets, to augment the training dataset.
Our approach, named Condor (literally meaning an eagle) not only searches for prey in the sky (metaphorically representing the general semantics of code) but also dives from the sky to capture prey on the ground (metaphorically representing the details of code). This hunting activity metaphorically illustrates how Condor (Code and its Nuanced Integrated Discriminator) can focus on both the overall structure and the fine-grained differences in code, thereby enhancing the performance of the discriminator.
2.2. Embedding-Level Contrast on Code Details
We employ a contrastive learning (Bui et al., 2021; Chen et al., 2024) strategy to make the model aware that textually similar yet functionally different code snippets should have distinct representations at the embedding level. Contrastive learning, known for enhancing the similarity between embeddings within the same category while reducing the similarity between those from different categories, is well-suited to our task. To fully utilize this mechanism, we design a similarity-based data sampling strategy tailored to code discrimination. By incorporating this design, we leverage contrastive learning to effectively capture subtle code differences, transforming them into meaningful parameter updates to refine the model’s code embedding representations.
2.2.1. Similarity-based Sampling
In contrastive learning, we focus on comparing samples with a high degree of textual similarity in the data sampling process, enabling the discriminator to differentiate between textually similar but functionally distinct code snippets. Each data sample in contrastive learning is represented as $\langle p,c_{a},c_{b},label_{a,b}\rangle$, where $c_{a}$ and $c_{b}$ are two textually similar code snippets, and $label_{a,b}$ indicates the relationship between them. Each contrastive learning sample takes into account both the problem description and the code representation, ensuring that the code embeddings are compared within the context of a specific problem. Based on whether the code can pass test cases (C: correct code that passes all test cases, E: error code that fails some test cases), the relationship between $c_{a}$ and $c_{b}$ can be classified into four categories: C-C, C-E, E-E, and E-C. Due to the symmetry between E-C and C-E in classification, we treat them as a single category. Since we primarily focus on the subtle differences between correct and incorrect code, we emphasize comparisons between C-C and E-C pairs. This helps the model recognize which code details affect functionality and which do not. Thus, the labels $label_{a,b}$ are divided into two key types: C-C (i.e., pairs with the same functionality) and E-C (pairs with different functionality). We assign a label of 1 to C-C pairs and 0 to E-C pairs and use these labeled data to train the supervised model for contrastive learning.
2.2.2. Contrastive Learning
Contrastive learning aims to make the embeddings of the same functionality pairs closer and those of different functionality pairs farther apart. This helps the model differentiate between correct code (C) and error code (E) at the representation level.
To achieve this goal, we first map the code and its associated problem representation into points in a two-dimensional space. We then quantify the similarity between code snippets by calculating the Euclidean distance between these points. We design a loss function based on Euclidean distance, aiming to maximize the distance between functionally correct and incorrect code pairs while minimizing the distance between functionally correct code pairs. The contrastive learning loss function is defined as follows:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left[label_{a,b}\cdot d(x_{a}^{i},x_{b}^{i})^{2}+(1-label_{a,b})\cdot\max(0,m-d(x_{a}^{i},x_{b}^{i}))^{2}\right],$
where $N$ is the total number of training samples, $label_{a,b}$ is a binary label for each pair, and $label_{a,b}=1$ indicates the same functionality pairs (P-P) and $label_{a,b}=0$ indicates different functionality pairs (N-P). $d(x_{a}^{i},x_{b}^{i})$ is the Euclidean distance between the two embeddings $x_{a}^{i}$ and $x_{b}^{i}$. $m$ is the margin that ensures dissimilar pairs are separated by at least a certain distance.
This loss encourages the same functionality pairs to have smaller distances while different functionality pairs are pushed apart by at least the margin $m$. For example, in Figure 2, this loss function encourages the model to increase the distance $dis_{1}$ (between the embeddings of correct and buggy code pairs) and decrease the distance $dis_{2}$ (between the embeddings of correct pairs). This optimization helps the model learn to distinguish between functionally correct and incorrect code, even when the text is similar.
2.3. Data-Level Augmentation on Code Details
In real-world development, programmers often make multiple changes while attempting to fix a bug, but not all intermediate code versions are recorded. For example, as shown in Figure 1, the user submitted two versions of the code—an incorrect version (on the left) and a corrected version (on the right). It can be observed that transitioning from the incorrect to the correct code requires two-step modifications, and simply fixing one part would not have resulted in a proper solution. For example, a potentially unrecorded version that combines the first two lines of the correct code (Submission 2) with the last three lines of the error code (Submission 1) would still be incorrect.
2.3.1. Intermediate Code Details
To ensure the model to recognize that the code must be correct in all details, we explore the potential paths from incorrect to correct code by incorporating partially fixed intermediate versions. We leverage these intermediate bug-fix data (e.g., as shown in the Figure 1, step 1 is modified in line 2 while step 2 remains unchanged in line 3) samples to enhance the training set. By introducing these partially fixed versions, the model gains a deeper understanding of the nuanced differences in code modifications, improving its ability to perceive and discriminate code correctness. This strategy effectively simulates the real bug-fixing process, helping the model know whether the current modifications are sufficient to correct the code.
To describe this process more clearly, we use the following notation to represent the various stages of modification: Let $E_{1}$ represent the initial incorrect code and $C_{1}$ represent the final correct code. We first compute $\text{diff}(E_{1},C_{1})$, which represents all the modifications required to transform the incorrect code $E_{1}$ into the correct code $C_{1}$. Next, we decompose $\text{diff}(E_{1},C_{1})$ into $n$ independent modification steps:
$\text{diff}(E_{1},C_{1})=\{\Delta_{1},\Delta_{2},\dots,\Delta_{n}\},$
Each $\Delta_{i}$ represents a part of the modifications needed to transform $E_{1}$ into $C_{1}$. Specifically, each $\Delta_{i}$ is represented as a diff hunk, which may consist of multiple consecutive lines of changes. If the difference between the two code versions involves only one hunk, the lines within that hunk are further split, treating each line as an individual modification. We then apply these modifications sequentially to $E_{1}$, generating a series of intermediate code versions $IE_{11},IE_{12},\dots,IE_{1n}$:
$IE_{11}=E_{1}+\Delta_{1}$
$IE_{12}=IE_{11}+\Delta_{2}$
$\vdots$
$C_{1}=IE_{1(n-1)}+\Delta_{n}$
Through this process, we generate multiple intermediate versions of the code as it progresses from incorrect to correct. Each intermediate version $IE_{1i}$ represents the code state after a partial modification. After generating all the intermediate code versions $IE_{11},IE_{12},\dots,IE_{1n}$, we further perform deduplication against the original training data to ensure no duplicate code snippets are introduced. Once deduplication is complete, these filtered intermediate code versions are added to the original training set, creating an augmented dataset. This approach allows the model to learn the subtle changes from partial to full fixes, improving its ability to understand and handle code details. This approach generates more training data with finer granularity and provides the model with multi-level information about the code repair process, helping the model better capture the progressive transformations from incorrect to correct code.
2.3.2. Discriminating
As a code discriminator, Condor intends to select functionally correct code from a set of candidate code snippets. Based on the embedding representations enriched with code details through contrastive learning, and the augmented training dataset containing intermediate code details, we train a binary discriminative model to determine whether a given code is functionally correct. Let the candidate code set be $\{C_{1},C_{2},\dots,C_{k}\}$.
During training, we first define the label $y_{i}\in\{0,1\}$, where $y_{i}=1$ indicates that the code $C_{i}$ is functionally correct, and $y_{i}=0$ indicates that the code is functionally incorrect. Then, we construct pairs $\langle C_{i},y_{i}\rangle$ to train a binary classification model as the discriminator. The cross-entropy loss function is applied to optimize the model, improving its ability to distinguish between correct and incorrect code. The loss function is defined as:
$\mathcal{L}=-\frac{1}{k}\sum_{i=1}^{k}\left(y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right),$
where $p_{i}=\sigma(f(\mathbf{V}_{i}))$ represents the likelihood that code $C_{i}$ is functionally correct, $f(\mathbf{V}_{i})$ is the logits value output by the model, and $\sigma(\cdot)$ is the sigmoid activation function. Specifically, $p_{i}$ is computed by applying the sigmoid function to the embedding logits, which maps the output to a probability range between 0 and 1, indicating the confidence of the model in the functional correctness of $C_{i}$.
During inference, given a set of candidate code snippets $\{C_{1},C_{2},\dots,C_{k}\}$, the model outputs the probability $p_{i}$ that each code snippet $C_{i}$ is functionally correct. We then select the candidate code with the highest probability as the most likely correct code:
$C_{\text{selected}}=\arg\max_{i}p_{i}$
2.4. CodeNanoFix Dataset
To better train a discriminator capable of detecting subtle code differences, it is essential to use data samples from real-world programming scenarios where programmers have written textually similar but functionally distinct code. Existing code generation datasets (Chen et al., 2021b; Austin et al., 2021; Hendrycks et al., 2021) are less suitable because these datasets lack real-world fine-grained modifications, and their training set relies on code generated by specific LLMs, making the discriminator’s effectiveness dependent on the particular LLM used. Additionally, identical functionality can be implemented in various ways, resulting in generated code that is not always textually similar. Therefore, we construct a new repair-related dataset, CodeNanoFix, comprising training, validation, and test sets to support the training of discriminators capable of detecting subtle changes in code.
2.4.1. Data Collection
CodeNanoFix consists of tuples of problem description, buggy code, and correct code, designed to evaluate human-written code with subtle differences. For example, Figure 1 illustrates two code snippets that are nearly identical in text but differ in functionality: Submission 1 on the left is incorrect (not passing the test cases), while submission 1 on the right is a correct solution corresponding to the problem description (passing the test cases). We collect the data samples from CodeNet (Puri et al., 2021), a large-scale collection of over 14 million code samples across over 50 programming languages, focusing on 4,000 problem-solving tasks from competitive programming. Each problem $p$ in the CodeNet dataset includes multiple submissions $S=(s_{1},s_{2},...)$ from different users and these submissions consist of both buggy and correct code snippets. Figure 1 depicts a user’s progression from an initial failed submission to a correctly revised version. We construct CodeNanoFix targeting examples where two submissions are textually similar but functionally different, specifically focusing on Python, the predominant language for evaluating code LLMs.
2.4.2. Data Cleaning
Data quality is crucial when building a code-related dataset. We observe several issues in the original CodeNet dataset that could potentially impact model training and evaluation. First, some code snippets in the dataset are mislabeled, such as code that fails certain test cases is incorrectly marked as functionally correct. To ensure label accuracy, we rerun all test cases and verify each label. We remove code samples where test results did not match the original labels, retaining only those where the test outputs are consistent with the original labels. Thus, we aim to maintain data consistency without introducing new labels, thereby avoiding potential label errors due to variations in test environments. Additionally, the problem descriptions in the original dataset are in different languages, which may interfere with the model’s understanding of the task. To address this, we use ChatGPT (OpenAI, 2024) to translate all problem descriptions into English for uniformity. These improvements enhance the dataset’s consistency and reliability, allowing it to reflect model performance in code discrimination tasks more accurately.
Table 1. The statistics of the CodeNanoFix dataset.
#Problem #Sample
#Average Tokens.
Problem
#Average Tokens.
Buggy Code
#Average Tokens.
Correct Code
#Average
Edit Distance
#Average Relative
Edit Distance
Training Set 480 80,906 187.64 109.65 112.82 15.52 0.11
Validation Set 565 4,742 202.31 198.52 273.73 22.78 0.10
Test set 120 3,583 178.68 129.24 133.36 19.86 0.12
2.4.3. Dataset Construction
After determining the problem descriptions and the labels corresponding to individual code snippets, we construct data samples in the format $\langle p_{i},c^{neg}_{i},c^{pos}_{i}\rangle$, where $c^{neg}_{i}$ represents an incorrect code submission for problem $p_{i}$, and $c^{pos}_{i}$ represents a correct submission. For a complex problem $p_{i}$, a user often needs multiple attempts to reach the correct solution. For example, with three submissions $S=(s1,s2,s3)$, $s3$ may be the correct code, while $s1$ and $s2$ contain bugs. To ensure that only minimal modifications are required for the model to fix buggy code, we calculate the Jaccard similarity (Jaccard, 1901) between the correct submission $s3$ and the buggy submissions $s1$ and $s2$, denoted as $j(s1,s3)$ and $j(s2,s3)$. We retain only those samples where the Jaccard similarity is greater than 0.9. Given the large number of users, and to ensure diversity in the constructed data samples, we group multiple users into one user group and follow the same process to create data samples. Table 1 shows the basic statistics of the CodeNanoFix dataset. The final dataset consists of training, validation, and test sets, each containing 480, 565, and 120 algorithmic problems and 80,906, 4,742, and 3,583 data samples. The validation set contains algorithmic problems that are generally more specialized, leading to a greater number of problems but fewer data samples. To avoid data leakage, we ensure that problem descriptions in the training and test sets are as distinct as possible, reducing any potential influence of training data on test data. This setup allows for effective evaluation of the model’s performance during training.
2.4.4. Dataset Metric
To verify data samples in the CodeNanoFix dataset requiring only minimal changes to fix bugs, we utilize the relative edit distance (RED) metric to quantify the similarity between buggy code and correct code. We first tokenize the code using the DeepSeek-Coder (Guo et al., 2024) tokenizer and calculate the number of tokens per code sample. Then, we compute the average edit distance (ED) between buggy and correct code versions to quantify the number of modification operations required for repairs. The results indicate that the average edit distance involves significantly fewer tokens than the total token count of the code itself. The RED metric is in the last column of the table, representing the ratio of required edit operations to the original token count (a value between 0 and 1). The specific formula is as follows:
RED=EDAVG(B_Tokens, $\text{RED}=\frac{\text{ED}}{\text{AVG(B\_Tokens, C\_Tokens)}},$
where $AVG$ indicates the average value of two numbers, $B\_Tokens$ represents the number of tokens in the buggy code, and $C\_Tokens$ represents the number of tokens in the correct code. The results show a typical RED of around 0.1, confirming that only minor adjustments are needed to change one code snippet to the other in the CodeNanoFix dataset. As a result, CodeNanoFix is more streamlined and consistent, making it better suitable for evaluating the discriminator’s ability to distinguish subtle changes in code.
3. Experimental Setup
In this section, we first present our research questions (RQs). Then, we present the benchmark datasets, baseline models, evaluation metrics, and implementation details used to evaluate Condor.
3.1. Research Questions
To evaluate the effectiveness of Condor, we ask the following questions:
RQ1: How well does Condor recognize subtle code differences and distinguish between textually similar but functionally different code?
To answer this question, we apply Condor to the CodeNanoFix benchmark dataset and evaluate its performance in classification and discrimination scenarios.
RQ2: Can Condor generalize to other code generation tasks?
To explore this, we evaluate its performance on common code generation datasets, comparing the Pass@1 performance before and after discrimination.
RQ3: How do the proposed code detail capture strategies affect model performance?
To address this, we conduct ablation experiments comparing the performance of the embedding-level contrastive learning strategy and the data-level intermediate code augmentation strategy.
3.2. Datasets
We introduce the CodeNanoFix benchmark dataset, which contains 80,906 training samples, 4,742 validation samples, and 3,583 test samples, specifically designed to assess a model’s ability to distinguish subtle code differences. This dataset focuses on code examples where only minor changes are needed for bug fixes, aiming to evaluate whether models can accurately capture code detail variations and effectively differentiate between functionally correct and incorrect code.
We conduct experiments on the APPS and MBPP benchmark datasets to further validate Condor’s generalization capabilities. The APPS dataset consists of 5,000 training problems and 5,000 test problems, covering a wide range of complex programming problems. We evaluate each baseline’s original code generation accuracy and its post-process accuracy on the test set, analyzing how Condor performs in more challenging code generation tasks. The MBPP dataset contains 374 training problems, 90 validation problems, and 500 test problems, primarily used for code generation evaluation. Through experiments on the MBPP dataset, we test Condor’s generalization ability and discrimination precision across problems of different scales.
By evaluating Condor on the CodeNanoFix, APPS, and MBPP datasets, we verify the model’s performance across various tasks and data scales, especially its ability to capture code details and generalize to diverse code generation tasks.
3.3. Baselines
We categorize the baselines into two types: baselines for training various discriminators and baselines for code generation. To make the discriminator more accessible, the models involved in training discriminators are typically smaller in parameter size compared to code generation models. Therefore, we select a random strategy, along with models based on CodeRanker, CodeT5, and DeepSeek-Coder-1.3B, as baselines for evaluating the trained discriminator’s performance. To provide a convincing comparison for our discriminator’s performance, we use high-performing models in software engineering tasks, specifically DeepSeek-Coder-Instruct-6.7B, Meta-Llama-3.1-Instruct-8B, and Meta-Llama-3.1-Instruct-70B, as baselines for code generation.
3.3.1. Baselines for Training Discriminators
For the discrimination-related tasks, we select simple strategies and models with smaller parameter sizes as baselines. The random strategy serves as a simple baseline by randomly selecting an output from the model’s generated results as the final output. This approach helps assess the variability of experimental results and ensures the robustness of the benchmark. CodeRanker (Inala et al., 2022) fine-tunes CodeBERT (Feng et al., 2020), an encoder model based on the BERT (Devlin et al., 2018) architecture, for the task of discriminating the outputs of LLMs. Trained with a masked language model on code data, CodeBERT excels in code representation learning, which enhances its applicability to various code understanding and classification tasks. CodeT5 (Wang et al., 2021) is an encoder-decoder structure that incorporates identifier-aware techniques to enhance code comprehension. Following existing studies (Alrashedy et al., 2024; An et al., 2024), we use the encoder of CodeT5 as a baseline for training the discriminator. DeepSeek-Coder-1.3B (Guo et al., 2024) is a decoder model based on the GPT architecture, trained on a large, high-quality code dataset, and performs exceptionally well in various code generation tasks, making it a solid baseline for training discriminators. To ensure that all baselines are compared at a high initial performance level, we generate training data for the MBPP and APPS datasets using code-generation baselines to serve as training samples for the discriminator. This approach not only provides high-quality training examples for the discriminator but also ensures consistency in the training data across all discriminators, facilitating a fair comparison.
3.3.2. Baselines for Code Generation
For code generation and repair tasks, we select high-performing, large-scale GPT-style models as baselines. DeepSeek-Coder-Instruct-6.7B (Guo et al., 2024) offers stronger code understanding and generation capabilities than its 1.3B version, making it a high-performance baseline for generation. Meta-Llama-3.1-Instruct-8B (Dubey et al., 2024), with 8 billion parameters, is a powerful code generation model based on the LLaMA architecture (Touvron et al., 2023; Roziere et al., 2023), specifically optimized for complex code generation scenarios. Meta-Llama-3.1-Instruct-70B, with 70 billion parameters, handles even greater code generation complexity, establishing itself as a top-performing baseline for code generation tasks.
3.4. Metrics
To comprehensively evaluate the performance of the Condor model in code classification and discrimination tasks, we use multiple evaluation metrics, including Precision, Recall, F1-score, and Pass@1 (Powers, 2020) to measure the model’s performance before and after discriminating.
3.4.1. Precision
Precision measures the proportion of correctly identified functional code among all code that the model predicted as correct. It is defined as:
$\text{Precision}=\frac{TP}{TP+FP}$
where $TP$ represents the number of correctly predicted functional code samples, and $FP$ represents the number of incorrectly predicted functional code samples. Precision reflects the model’s performance in minimizing false positives.
3.4.2. Recall
Recall measures the proportion of all actual correct code that the model successfully identified. It is defined as:
$\text{Recall}=\frac{TP}{TP+FN}$
where $FN$ represents the number of actual correct code samples that the model incorrectly classified as incorrect. Recall highlights the model’s ability to identify all correct code samples.
3.4.3. F1-score
The F1-score is the harmonic mean of precision and recall, combining both metrics to provide a balanced evaluation. It is defined as:
[…]
The F1-Score is useful when precision and recall need to be balanced, offering a more comprehensive reflection of the model’s overall performance.
3.4.4. Pass@1
Pass@1 is an important metric in code generation tasks, representing the accuracy of the top-1 candidate code (the model’s first choice). It is primarily used to assess the accuracy of code generation and is defined as:
[…]
When evaluating the model’s performance before and after discrimination, we calculate both the original Pass@1 from the generated code and the Pass@1 after applying the discrimination mechanism, comparing the performance improvement.
By combining Precision, Recall, F1-score, and Pass@1, we can comprehensively evaluate the Condor model’s performance in code classification and discrimination tasks. This approach captures the model’s ability to identify correct code while also assessing improvements brought by the discrimination mechanism across different evaluation stages.
3.5. Implementation Details
We train two versions of Condor with different parameter scales (i.e., Condor-110M and Condor-1.3B) based on these strategies. We choose to use CodeT5’s encoder and DeepSeek-Coder-1.3B as the base models to train Condor at different parameter scales for two main reasons. First, both models demonstrate strong performance in various software engineering tasks and are widely used in the field, providing a solid foundation. Second, compared to larger models like DeepSeekCoder-Instruct-6.7B or Code-Llama-3.1-Instruct-70B, they are relatively small, which helps us control the parameter size of the discriminator, making training and deployment more efficient across a variety of tasks. The margin value $m$ for the contrastive learning loss function is set to 2 to reflect the typical magnitude of semantic differences between correct and incorrect code.
We conduct our experiments on a machine running the Ubuntu operating system, equipped with 8 NVIDIA A100 GPUs, each with 80GB of memory, and 500GB of system memory. This hardware setup provides the necessary support for large-scale model training and dataset loading. To ensure the reproducibility of our experiments, we implement the models and experiments using the Transformers and Huggingface (Huggingface, 2024b) libraries. Additionally, to accelerate the training process, we utilize DeepSpeed (Rajbhandari et al., 2020) and Accelerate (Huggingface, 2024a) libraries, which optimize computational efficiency in a multi-GPU environment, significantly reducing training time.
4. Results
4.1. RQ1: Effectiveness of Condor
We evaluate Condor’s effectiveness from two perspectives. First, we assess whether it can distinguish subtle code differences by treating the CodeNanoFix dataset as a classification task, determining whether the model can effectively classify between correct and incorrect code. Second, we use code LLMs to repair buggy code in the CodeNanoFix dataset and have Condor discriminate the generated outputs to see whether it enhances model performance after discrimination.
4.1.1. Effectiveness on Classification Scenario
Table 2. Condor’s Performance in Classification.
Model Precision Recall F1-Score
CodeRanker (110M) 59.42 77.56 67.29
CodeT5 (110M) 60.09 76.95 67.48
DeepSeek-Coder (1.3B) 71.99 62.77 67.06
Condor (110M) 65.99 80.04 72.34
Condor (1.3B) 74.39 72.40 73.38
Table 2 shows the classification performance of the models with different parameter sizes on the CodeNanoFix dataset. From the table, it can be observed that smaller models (such as CodeRanker and CodeT5) have lower precision but higher recall, while larger models (such as DeepSeek-Code) demonstrate higher precision but lower recall, achieving more balanced overall performance.
Two factors can explain these results. First, smaller models, with their limited numbers of parameters, struggle to distinguish subtle code differences accurately. As a result, they tend to broadly label potentially correct code to capture as many functionally correct samples as possible, leading to higher recall. However, this broad labeling strategy also increases the likelihood of false positives, resulting in lower precision. For example, CodeRanker and CodeT5 have recall values of 77.56% and 76.95%, respectively, but their precision only reaches 59.42% and 60.09%. Second, larger models, with more parameters, can capture finer-grained code features, making them more precise in identifying correct code and thus improving precision. However, as these models become more conservative, they are more cautious when labeling code as correct, potentially missing some functionally correct but borderline samples, leading to lower recall. For example, DeepSeek-Coder (1.3B) achieves a precision of 71.99% but a recall of only 62.77%.
Condor introduces embedding-level and data-level strategies to capture the code details, enhancing the model’s sensitivity to functional differences in code, and significantly improving the overall performance of models of various sizes. These strategies enable Condor to better capture subtle code changes, improving precision and recall for both small and large models. Specifically, Condor (110M) boosts precision to 65.99% and recall to 80.04%, showing a substantial improvement over CodeRanker and CodeT5. Meanwhile, Condor (1.3B) increases precision to 74.39% and recall to 72.40%, achieving a more balanced performance, with an F1-Score of 73.38%, higher than all other models. Overall, regardless of model size, Condor’s code-detail-aware strategies achieve significant improvements in both precision and recall, resulting in enhanced overall performance.
4.1.2. Effectiveness on Discrimination Scenario
Table 3. Condor’s discriminative performance in the CodeNanoFix dataset.
Model’s Pass@1
Deepseek-Coder
Instruct (6.7B)
Meta-Llama-3.1
Instruct (8B)
Meta-Llama-3.1
Instruct (70B)
Original 43.23 36.81 52.64
Random 43.23 36.75 53.39
CodeRanker (110M) 42.84 37.48 57.69
CodeT5 (110M) 42.19 37.65 56.93
DeepSeek-Coder (1.3B) 43.54 37.32 58.90
Condor (110M) 43.73 39.60 59.64
Condor (1.3B) 50.85 45.27 62.63
Table 3 presents the performance of various discriminators on the CodeNanoFix dataset, evaluating the functional correctness of the top-1 code snippet using the Pass@1 metric after selecting outputs from the LLMs. The performance of the original generation models varies significantly, with Deepseek-Coder Instruct 6.7B achieving a Pass@1 of 43.23%, Meta-Llama-3.1 Instruct 8B of 36.81%, and Meta-Llama-3.1 Instruct 70B performing the best of 52.64%.
After applying a discriminator to the generated outputs, we observe that Condor (1.3B) significantly outperforms other discriminators. For example, Condor (1.3B) raises the Pass@1 score on Deepseek-Coder Instruct (6.7B) to 50.85%, resulting in a 17.63% increase compared to the original result. On Meta-Llama-3.1-Instruct (70B), Pass@1 reaches 62.63%, reflecting an improvement of nearly 19% over the original result. While Condor (110M) also shows improvements across multiple models, the gains are less pronounced compared to their larger counterparts, highlighting the importance of parameter size in capturing complex code details. In contrast, other discriminators (i.e., CodeRanker and CodeT5) also enhance the quality of generated results but yield smaller improvements, particularly in more complex generation tasks. The results of random selection show little difference from the original models, indicating that random selection has minimal impact on model performance.
These results demonstrate that Condor’s incorporation of a code detail representation significantly enhances its ability to capture subtle code differences. Whether using a small or large model, Condor substantially improves the quality of selected generated code, demonstrating superior discrimination performance.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ1: The experimental results demonstrate that code details play a crucial role in improving both classification and discrimination performance. By introducing code detail-aware strategies, Condor significantly enhances its ability to capture subtle code differences, leading to improved precision and recall across different models. For example, Condor (1.3B) achieves a precision of 74.39% and a recall of 72.40%. In the discrimination scenario, Condor’s performance is particularly outstanding. On Meta-Llama-3.1 Instruct 70B, Condor (1.3B) increases Pass@1 from 52.64% to 62.63%. This demonstrates Condor’s ability to accurately identify correct code and exhibit strong discriminative capabilities across models of various sizes.
4.2. RQ2: Generalizability to Code Generation Tasks
Tables 4 and 5 present the discrimination capabilities of our Condor model on the APPS and MBPP datasets, using the Pass@1 metric to measure Condor’s effectiveness in selecting LLM outputs. From these experimental results, we can observe Condor’s generalization ability in code generation tasks.
Table 4. Condor’s generalization capability in APPS.
Model’s Pass@1
Deepseek-Coder
Instruct (6.7B)
Meta-Llama-3.1
Instruct (8B)
Meta-Llama-3.1
Instruct (70B)
Original 9.40 6.06 10.16
Random 9.60 6.18 11.74
CodeRanker (110M) 11.56 9.22 21.30
CodeT5 (110M) 12.18 9.36 21.62
DeepSeek-Coder (1.3B) 12.10 9.64 22.42
Condor (110M) 12.24 10.32 22.32
Condor (1.3B) 14.68 12.08 25.10
Table 5. Condor’s generalization capability in MBPP.
Model’s Pass@1
Deepseek-Coder
Instruct (6.7B)
Meta-Llama-3.1
Instruct (8B)
Meta-Llama-3.1
Instruct (70B)
Original 59.00 54.00 71.40
Random 59.60 55.20 71.60
CodeRanker (110M) 63.20 56.60 72.20
CodeT5 (110M) 61.60 58.20 72.80
DeepSeek-Coder (1.3B) 64.20 57.60 73.60
Condor (110M) 63.20 60.40 73.00
Condor (1.3B) 65.00 59.00 75.20
In the APPS dataset (Table 4), due to the challenging nature of the algorithmic problems, the original generation models show relatively low performance. Deepseek-Coder Instruct 6.7B achieves a Pass@1 of 9.40%, Meta-Llama-3.1 Instruct 8B reaches 6.06%, and the Meta-Llama-3.1 Instruct 70B version achieves 10.16%. The results of random selection differ little from the original outputs, for example, on Deepseek-Coder Instruct 6.7B, Pass@1 only increases from 9.40% to 9.60%. In contrast, traditional discriminators like CodeRanker (110M) and CodeT5 (110M) show more significant improvements, with CodeRanker (110M) boosting Meta-Llama-3.1 Instruct 70B’s Pass@1 to 21.30%, and CodeT5 (110M) slightly higher at 21.62%. As the model size increases, DeepSeek-Coder (1.3B) further improves Pass@1 to 22.42%, demonstrating solid discriminative ability. Condor, enhanced by its code detail-awareness strategies, performs exceptionally well on the APPS dataset. Condor (110M) surpasses CodeRanker and CodeT5, increasing Meta-Llama-3.1 Instruct 8B’s Pass@1 to 10.32%, showcasing its discrimination advantage even in smaller models. Condor (1.3B) performs the best, achieving the highest Pass@1 across all models. For example, on the Meta-Llama-3.1-Instruct (70B), it boosts Pass@1 to 25.10%, representing an improvement over 147% from the original 10.16%. This demonstrates Condor’s strong ability to capture code details and significantly enhance performance in discrimination tasks.
The experiments on the MBPP dataset (Table 5) further validate Condor’s generalization ability. Due to the relatively simpler nature of the MBPP problems, the original generation models’ Pass@1 scores are higher across the board. For example, Deepseek-Coder Instruct 6.7B achieves a Pass@1 of 59.00%, Meta-Llama-3.1 Instruct 8B reaches 54.00%, and Meta-Llama-3.1 Instruct 70B reaches 71.40%. Similar to the APPS dataset, random selection yields almost no improvement. CodeRanker (110M) and CodeT5 (110M) show more notable gains in discrimination performance, with CodeT5 (110M) improving Meta-Llama-3.1 Instruct 70B’s Pass@1 to 72.80%, slightly higher than CodeRanker (110M) at 72.20%. DeepSeek-Coder (1.3B) performs steadily on this dataset, raising Pass@1 to 73.60%. Similarly, Condor performs outstandingly on the MBPP dataset, especially Condor (1.3B), which further improves Meta-Llama-3.1 Instruct 70B’s Pass@1 to 75.20%, the best result. Condor (110M) also performs well on smaller models, boosting Meta-Llama-3.1 Instruct 8B’s Pass@1 to 60.40%, surpassing DeepSeek-Coder and other discriminators. Overall, Condor demonstrates competitiveness not only in smaller models but also in larger models, well utilizing its ability to capture code details and significantly improving Pass@1 accuracy.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ2: The experimental results demonstrate that Condor exhibits outstanding generalization capabilities, effectively improving discrimination accuracy on both the challenging APPS dataset and the relatively simpler MBPP dataset. On the APPS dataset, Condor (1.3B) increased Meta-Llama-3.1 Instruct 70B’s Pass@1 to 25.10%, compared to only 10.16% from the original generation. Similarly, on the MBPP dataset, Condor also performs exceptionally well, with Condor (1.3B) raising Meta-Llama-3.1 Instruct 70B’s Pass@1 to 75.20%, surpassing the original 71.40%. This indicates that Condor consistently enhances Pass@1 performance across tasks of various difficulty, showcasing its strong ability to capture subtle code details.
4.3. RQ3: Ablation Study
Table 6. Condor’s Ablation Study. The ‘w/o’ indicates without a specific strategy.
Model Precision Recall F1-Score
Pass@1 of DeepSeek-
Coder-Instruct (6.7B)
Pass@1 of Meta-Llama
-3.1-Instruct (8B)
Pass@1 of Meta-Llama
-3.1-Instruct (70B)
Condor (1.3B) 74.39 72.40 73.38 50.85 45.27 62.63
w/o Contrastive Learning 73.12 72.06 72.59 48.03 42.37 60.95
w/o Intermediate Data 67.98 74.66 71.16 45.29 38.54 59.84
Table 6 presents the ablation study results for our Condor model, analyzing performance changes on the CodeNanoFix dataset between the full model and its variants with specific strategies removed. Overall, the full Condor (1.3B) model delivers the best results across all metrics, with a Precision of 74.39%, Recall of 72.40%, and F1-score of 73.38%. It also excels in discrimination tasks, achieving a Pass@1 of 62.63% on Meta-Llama-3.1-Instruct (70B), highlighting its outstanding ability to capture code details and improve Pass@1 accuracy.
When the embedding-level contrastive learning strategy is removed, precision drops to 73.12%, and recall also decreases to 72.06%, resulting in a lower F1-score of 72.59%. This indicates that contrastive learning plays a crucial role in enhancing the model’s ability to capture functional code differences. The discrimination performance also declines, with Pass@1 on Meta-Llama-3.1-Instruct (70B) dropping from 62.63% to 60.95%, demonstrating the significant contribution of contrastive learning to discrimination tasks. When the data-level intermediate data augmentation strategy is removed, the model shows more pronounced changes. Although recall increases slightly to 74.66%, suggesting the model becomes more conservative in recognizing correct code, precision drops significantly to 67.98%, leading to an overall F1-score decrease to 71.16%. The discriminative performance also declines, with Pass@1 on Meta-Llama-3.1-Instruct (70B) falling to 59.84%. This demonstrates the critical role of the intermediate data strategy in improving the model’s overall performance.
Figure 3. Illustration of the impact of contrastive learning on code representations. The first three subplots show the 2D embeddings of code after training for 1 epoch (top left), 5 epochs (top right), and 40 epochs (bottom left). The bottom right subplot illustrates the changes in the average distance between correct code snippets and between error and correct code snippets as the number of training epochs increases.
Figure 3 visually demonstrates the impact of contrastive learning on model embedding representations. The subplots in the top left, top right, and bottom left show the progression of code embeddings projected in 2D space as training epochs increase. These projections, recorded before calculating the loss during contrastive learning, reveal how the embeddings evolve and separate as the model learns to distinguish correct from incorrect code. From these subplots, we observe that as training deepens, the embeddings of correct code snippets gradually cluster closer to the center, while the embeddings of buggy code samples become more dispersed, moving farther from the center of the correct code. This indicates that contrastive learning effectively brings representations of correct code closer together while increasing the distance between correct and error code, enhancing the model’s ability to discriminate code correctness. The bottom right subplot shows the average distances between correct code pairs (C-C) and between error and correct code pairs (E-C). As training progresses, the C-C distance gradually decreases, while the E-C distance increases. This trend further demonstrates that the contrastive learning strategy effectively assists the model in distinguishing fine-grained differences between correct and incorrect code representations at the embedding level.
Leveraging both contrastive learning and intermediate data augmentation strategies to capture the code details, the full Condor (1.3B) model achieves a balance between high precision and recall in classification tasks and delivers optimal performance in discrimination tasks. When these strategies are removed, the model’s ability to capture code details and Pass@1 accuracy declines, confirming the positive impact of both strategies on Condor’s final performance.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ3: The ablation study results show that both the contrastive learning and intermediate data augmentation strategies effectively help the model capture detailed information in the code, significantly enhancing overall model performance. When contrastive learning is removed, both the model’s F1-score and its discrimination performance decline, indicating that contrastive learning is crucial for capturing functional code differences. Similarly, removing the intermediate data augmentation strategy slightly improves recall but significantly reduces precision by 8.6% and overall discrimination performance, highlighting this strategy’s crucial role in enhancing the model’s accuracy and discriminative ability.
5. Threats to Validity
Threats to internal validity might come from the implementation of our experiments. To mitigate this threat, we use open-source models available on HuggingFace (Huggingface, 2024b) to minimize potential implementation errors. Additionally, to ensure the reproducibility of our experiments, we fix the random seed, ensuring consistency across different experimental results and reducing interference from randomness.
Threats to external validity may arise from two factors. The first factor may come from the choice of the base models. To mitigate this risk, we select the widely used and representative models CodeT5 and DeepSeek-Coder. CodeT5 is recognized as a classic encoder-decoder framework that has been validated and applied across multiple tasks, demonstrating strong versatility. DeepSeek-Coder is a decoder model that performs exceptionally well in a variety of software engineering tasks, providing a solid and reliable foundation for our experiments with our Condor model. The second potential threat comes from the downstream datasets used in our experiments. To mitigate this risk, we select the widely used MBPP and APPS datasets. These datasets are extensively used to evaluate code generation tasks and are commonly used to assess the performance of LLMs, providing strong support for the robustness of our experimental results.
Threat to construct validity may arise from the choice of evaluation metrics. To mitigate this, we use widely accepted metrics: precision, recall, and F1-score for classification tasks to comprehensively assess the model’s discriminative ability. For discrimination tasks, we apply the Pass@1 metric, commonly used in the code generation domain, to evaluate the model’s post-selection performance. These metrics effectively measure the model’s performance across different tasks, ensuring comprehensive and reliable evaluations.
6. Related Work
6.1. Code Discriminators
Existing approaches that improve code generation performance by selecting the outputs of LLMs can be categorized into two types: execution-based approaches and non-execution-based approaches. Execution-based approaches require executing the code and then using the feedback from the execution as additional information to enhance the discriminative effectiveness. Among them, MBR-EXEC, CodeT (Chen et al., 2023; Shi et al., 2022) requires executing the code to select the LLM’s outputs. However, in real-world development scenarios, it is not always easy to obtain the necessary test cases or testing environments for execution. Moreover, executing code generated by LLMs often raises safety concerns. Consequently, execution-based approaches face significant challenges in terms of flexibility. For flexibility, we focus on the non-execution-based approaches. Coder-Reviewer reranking (Zhang et al., 2023) uses prompt engineering by feeding the generated code to the original model to generate descriptions. It then selects the code by calculating the similarity between the generated description and the original description. However, as the current code model is quite large, making it costly to use the original model to assist in selecting generated outputs, which falls short of the efficiency requirements for real-world code generation tasks. Additionally, this approach is highly sensitive to prompt design and the model’s ability to interpret natural language, especially when problem descriptions are complex (e.g., APPS dataset). This discrepancy in the complexity of generating problem descriptions versus code can significantly impact the accuracy and consistency of the generated results. CodeRanker (Inala et al., 2022), which is similar to our work, improves model accuracy by incorporating a smaller classification model. Specifically, CodeRanker fine-tunes the CodeBERT model for code classification tasks, which in turn enhances the discrimination of the model’s outputs. Unlike existing models that simply leverage the inherent capabilities of models to represent code, Condor employs contrastive learning and data augmentation techniques to better detect subtle code details, thereby enhancing the discrimination performance.
6.2. Datasets for Code Discrimination
Current code discriminators commonly use code generation datasets as evaluation benchmarks. For example, CodeRanker, CodeT (Inala et al., 2022; Chen et al., 2023) have been evaluated on APPS, MBPP, HumanEval (i.e., code generation datasets for Python programming language) (Hendrycks et al., 2021; Austin et al., 2021; Chen et al., 2021b), and MBR-EXEC, Coder-Reviewer reranking (Shi et al., 2022; Zhang et al., 2023) have been evaluated on Spider (i.e., generation dataset for SQL code), NL2Bash (i.e., generation dataset for Bash code ) (Yu et al., 2018; Lin et al., 2018). Although these code-generation datasets can be used to evaluate the overall performance of discriminators, they fall short in assessing the discriminators’s true potential to detect subtle code details. In code generation datasets, inputs are typically natural language descriptions prompting the model to generate corresponding code. For the same functionality, there may be multiple implementation approaches, so the generated code is not always textually similar. Additionally, these datasets rely on outputs generated by LLMs, making their effectiveness dependent on the specific LLM used. Therefore, we construct the CodeNanoFix dataset from real user-submitted code, focusing on detailed changes in the code repair process to provide a more direct and effective evaluation of the discriminator’s performance in distinguishing code details.
6.3. Code Repair Datasets
There are some studies focus on the construction of code repair datasets, such as Defects4J, Bugs.jar, Bugs2Fix, QuixBug, IntroClass, HumanEvalPack (Just et al., 2014; Saha et al., 2018; Lu et al., 2021; Lin et al., 2017; Le Goues et al., 2015; Muennighoff et al., 2023). However, these datasets are not designed to evaluate the discriminator’s ability to detect subtle code details. The reasons are as follows. (1) For repository-level repair datasets like Defects4J and Bugs.jar (Just et al., 2014; Saha et al., 2018), the process involves multiple stages, including defect localization, code repair, and patch detection. LLMs can hardly directly generate a series of code candidates to discriminate without specific fine-tuning or prompt design. In contrast, our CodeNanoFix dataset includes independent problem descriptions and code, allowing LLMs to directly generate fixed code for buggy code. Additionally, CodeNanoFix focuses on fine-grained code differences rather than extensive modifications, making it easier to evaluate the model’s potential. (2) The Bugs2Fix, QuixBug, IntroClass, and HumanEvalPack datasets are function-level repair datasets, but they are still not suitable for evaluating the discriminators. The Bugs2Fix (Lu et al., 2021) is a Java code repair dataset that includes erroneous Java code with variable anonymization. Due to the anonymization and lack of natural language descriptions of the code’s functionality, this dataset focuses more on syntactic and type errors rather than semantic errors. The QuixBug dataset (Lin et al., 2017) contains 40 erroneous Python code samples, with only a test set and no training set, aimed at evaluating code repair effectiveness. However, this dataset focuses on different repair problems and lacks fine-grained variations of code for the same repair task, making it challenging to evaluate the capabilities of discriminators. The IntroClass dataset (Le Goues et al., 2015) provides tens of thousands of erroneous C language answers for six programming tasks, while CodeNanoFix is sourced from thousands of different problems, offering greater semantic diversity. The HumanEvalPack (Muennighoff et al., 2023) dataset is based on 164 problems from the HumanEval (Chen et al., 2021b) dataset, with each problem containing a manually injected bug. However, these samples are created by intentionally introducing errors into correct code, rather than reflecting real-world bug-fixing scenarios, making it difficult to capture the true nature of code repair. Additionally, each problem in this dataset corresponds to only one bug and its corrected version, lacking coverage of diverse repair contexts for the same problem and thus failing to effectively showcase the variety of potential fixes. Therefore, we introduce a novel repair dataset, CodeNanoFix, that addresses the limitations of existing datasets in evaluating a model’s ability to recognize subtle code differences. In the CodeNanoFix dataset, the differences between the buggy and fixed code are minimal, enabling a more effective assessment of the model’s discriminative capability.
7. Conclusion
We propose Condor, a code discriminator that focuses on fine-grained code differences. To enhance the model’s ability to discriminate subtle changes, we propose an embedding-level code detail contrastive strategy and a data-level code detail augmentation strategy. Specifically, we first employ a contrastive learning approach to optimize the code representations of the original base model, making it more sensitive to changes in code details. Then, we extract intermediate modification data from the code revision process to further enrich the training data for the discriminator, enhancing its ability to discern code details. To evaluate the performance of the discriminator, we construct the CodeNanoFix dataset, which consists of buggy code resulting from minor differences and their corresponding correct versions. This dataset is used to assess the capabilities of various discriminators. Experimental results demonstrate that Condor outperforms other discriminators in classification performance and excels at selecting generated outputs from LLMs. Additionally, Condor is applied to discrimination tasks on other code-generation datasets, with experimental results demonstrating strong generalization capabilities, as indicated by a significantly higher Pass@1 metric compared to other models. Ablation studies further confirm that both strategies employed by Condor positively impact its overall performance.
8. Data Availability
Our code and dataset are available at the link: https://github.com/LIANGQINGYUAN/Condor